OpenStudy (anonymous):

What is the sum of all square numbers from 0 to 101?

OpenStudy (callisto):

we know that the smallest square number there is 1, largest is 100 (10^2) Since 1^2 +2^2 +3^2 +...+n^2 = (n)(n+1)(2n+1)/6 Therefore sum required = 1^2 +2^2 +3^2 +...+10^2 = (10)(10+1)(2x10+1)/6 = Can you work it out?

OpenStudy (radar):

100+81+64+49+36+25+16+9+4+1 = ------- Or use Callisto method, it is easier.